# Model Predictive Control via Probabilistic Inference: A Tutorial

### v1

@@ Abstract
Model Predictive Control (MPC) is a fundamental framework for optimizing robot behavior over a finite future horizon. While conventional numerical optimization methods can efficiently handle simple dynamics and cost structures, they often become intractable for the nonlinear or non-differentiable systems commonly encountered in robotics. This article provides a tutorial on probabilistic inference-based MPC, presenting a unified theoretical foundation and a comprehensive overview of representative methods. Probabilistic inference-based MPC approaches, such as Model Predictive Path Integral (MPPI) control, have gained significant attention by reinterpreting optimal control as a problem of probabilistic inference. Rather than relying on gradient-based numerical optimization, these methods estimate optimal control distributions through sampling-based techniques, accommodating arbitrary cost functions and dynamics. We first derive the optimal control distribution from the standard optimal control problem, elucidating its probabilistic interpretation and key characteristics. The widely used MPPI algorithm is then derived as a practical example, followed by discussions on prior and variational distribution design, tuning principles, and theoretical aspects. This article aims to serve as a systematic guide for researchers and practitioners seeking to understand, implement, and extend these methods in robotics and beyond.
@@
Model Predictive Control, Probabilistic Inference, Sampling-Based MPC, Model Predictive Path Integral Control, Model-Based Reinforcement Learning
@@ 1 Introduction
Establishing a universal control theory applicable to diverse robots and tasks has long been a fundamental goal in robotics research. For complex tasks, we often require control methods that optimize actions while predicting future states. Model Predictive Control (MPC) and related optimal control frameworks are built upon this principle. MPC formulates robot control and task execution as finite-horizon optimal control problems, solving them at each time step to determine sophisticated robot behaviors.
To apply MPC to a specific application, we design cost functions and system dynamics for the target robot and task. However, the resulting formulations often exhibit strong nonlinearity, making them intractable for numerical optimization methods. In many practical applications, what we truly need is not infinite-horizon stability or convergence guarantees, but rather an optimization framework capable of reliably solving a broad class of optimal control problems with straightforward implementation.
@@
This article provides a tutorial on the theory and applications of probabilistic inference-based approaches to solving MPC optimal control problems, a topic that has attracted robotic research attention in recent years. This framework belongs to the class of sampling-based MPC methods, which randomly sample robot actions and determine optimal behaviors from these samples. Consequently, it is fundamentally applicable to optimal control problems with arbitrary dynamics and cost functions.
Attempts to solve optimal control problems through probabilistic inference date back to the early 2000s, beginning with path integral control [kappen2005path], extending to information-theoretic frameworks through Model Predictive Path Integral Control (MPPI) [williams2018information], and reaching connections to Reinforcement Learning (RL) through the general framework of Control as Inference (CaI) [levine2018reinforcement]. This article begins with an exposition of general frameworks such as CaI, derives the widely-used MPPI algorithm, and introduces advanced methods and practical considerations for applications. In addition to the aforementioned references, this work draws upon [wang2021variational, abdolmaleki2018maximum].
2 Preliminaries
@@
Figure 1: Overview of probabilistic inference-based MPC. The framework consists of two stages: (1) deriving the optimal control distribution from the optimal control problem, and (2) generating control input sequences from the optimal control distribution.
2.1 Optimal Control Problem Formulation
@@
Here, $J$ denotes the cost function, $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ represent the state and control input at time $t$, $\mathbf{u}_{0:T-1}=[\mathbf{u}_{0},\dots,\mathbf{u}_{T-1}]$ is the control input sequence, $\mathbf{x}_{0:T}=[\mathbf{x}_{0},\dots,\mathbf{x}_{T}]$ is the state sequence, $\mathbf{f}$ represents the system dynamics, and $\mathbf{x}_{0}$ is the observable initial state. Note that most optimal control problems include constraints on states and control inputs. In probabilistic inference MPC, input constraints can be directly incorporated as part of the dynamics, while state constraints cannot be handled directly and are typically embedded in the cost function using indicator functions. For details, refer to the experiments in [williams2018information].
Users design the cost function $J$ and dynamics $\mathbf{f}$, solve Eq. (1) in real-time, and control the robot by applying $\mathbf{u}^{*}_{0:T-1}$. When $J$ and $\mathbf{f}$ have simple forms (e.g., quadratic), efficient solutions can be obtained using numerical optimization methods. However, when $J$ or $\mathbf{f}$ exhibit nonlinearity or lack analytical derivatives, solving via numerical optimization becomes extremely challenging.
@@ 2.2 Motivations for Probabilistic Inference-Based MPC
When numerical optimization proves infeasible, our remaining option is computationally intensive sampling-based methods. The most straightforward approach is random shooting method, i.e., randomly generate a large number of control input samples $\mathbf{u}_{0:T-1}$, simulate state trajectories using Eq. (1b), compute the cost $J$ for each trajectory via Eq. (1a), and select the control input sequence with the minimum cost. However, as shown in Fig. 2, this method suffers from extremely poor sample efficiency and faces the curse of dimensionality, where the required number of samples grows exponentially with input dimensionality.
Figure 2: A comparison of sampling-based MPC methods in a vehicle obstacle avoidance task [honda2024stein]. Left: Random shooting method. Right: Model Predictive Path Integral control (MPPI). Random shooting method finds a single optimal solution from random samples, leading to poor sample efficiency. In contrast, MPPI identifies an optimal control distribution, improving sample efficiency significantly.
To overcome the sample efficiency challenges, We therefore introduce a probabilistic inference framework. Specifically, as illustrated in Fig. 1, we approximate the probabilistic distribution of optimal control inputs (hereafter referred to as the optimal control distribution) using samples and generate control input sequences from this distribution. Instead of seeking a single optimal solution from random samples, identifying a distribution with inherent breadth based on samples significantly improves sample efficiency, as shown in Fig. 2. Beyond improved sample efficiency, formulating the problem as distribution optimization offers several additional advantages:
  • Enables representation of stochastic behavioral variability.
  • Consists primarily of parallel independent operations, facilitating straightforward parallelization on CPUs/GPUs.
  • Allows end-to-end differentiable implementation using libraries such as PyTorch.
In particular, stochastic behavioral variability is valuable for representing exploration behavior and plays a crucial role in model-based reinforcement learning [hansen2022temporal, okada2020variational].
3 Probabilistic Inference-Based MPC
In this section, we introduce the foundations of the probabilistic inference-based MPC framework. As shown in Fig. 1, the scheme involves two key steps: deriving the optimal control distribution from the optimal control problem in Eq. (1) and then generating control inputs to apply to the robot from this optimal control distribution. Thus, two fundamental questions arise:
@@ 3 Probabilistic Inference-Based MPC
3.1 Derivation of the Optimal Control Distribution
We begin by deriving the optimal control distribution from the optimal control problem in Eq. (1). First, we state the conclusion: under some assumptions, the probability density function $\pi^{*}$ of the optimal control distribution can be expressed by the following remarkably simple formula:
@@ 3.1 Derivation of the Optimal Control Distribution
Here, $Z$ is the normalization constant, $\lambda$ is the temperature parameter, $J_{\tau}(\mathbf{u}_{0:T-1})$ is the trajectory cost obtained from the control input $\mathbf{u}_{0:T-1}$ via Eq. (1), and $p(\mathbf{u}_{0:T-1})$ represents the prior distribution over control inputs. Thus, the optimal control distribution is expressed as the product of a Boltzmann distribution and a prior distribution [^1] A similar expression appears in Direct Preference Optimization [rafailov2023direct], known for alignment methods in large language models.. While we detail the Boltzmann distribution later, it is a distribution that expresses trajectory optimality. We now derive this expression in three steps.
3.1.1 Graphical Model Representation of the Optimal Control Problem
Figure 3: Graphical model representation of the optimal control problem.
We begin by representing the optimal control problem in Eq. (1) as a graphical model. A graphical model depicts dependencies between random variables, where nodes represent random variables and edges represent dependencies. Equation (1) can be represented by the graphical model shown in Fig. 3. As illustrated, $\mathbf{x}_{t+1}$ depends on $\mathbf{u}_{t}$ and $\mathbf{x}_{t}$, representing Eq. (1b). Additionally, edges from $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ lead to $\mathcal{O}_{t}$, called the optimality variable, i.e., a virtual binary variable that returns whether the state and input are optimal ($\mathcal{O}_{t}=1$) or not optimal ($\mathcal{O}_{t}=0$). This $\mathcal{O}_{t}$ represents the cost function in Eq. (1a).
Based on this graphical model, we derive the probability distribution of the optimal trajectory $\tau=[\mathbf{x}_{0:T},\mathbf{u}_{0:T-1}]$. Setting $\mathcal{O}_{0:T}=[\mathcal{O}_{0},\dots,\mathcal{O}_{T}]$, Bayes’ theorem gives the optimal trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ as:
@@ 3.1.1 Graphical Model Representation of the Optimal Control Problem
where $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$ is the optimality likelihood, representing the probability that $\tau$ is an optimal trajectory.
3.1.2 Approximating the Optimal Control Distribution via Kullback-Leibler (KL) Divergence Minimization
While we modeled the optimal trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ in the previous section Eq. (3), the virtual binary variable $\mathcal{O}_{0:T}$ cannot be computed in advance, making this distribution intractable. We therefore employ variational inference to approximate the optimal trajectory distribution. Specifically, we introduce a distribution $\pi(\tau)$ that does not depend on $\mathcal{O}_{0:T}$ and minimize the KL divergence $\mathbb{D}_{\rm{KL}}$ between $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ and $\pi(\tau)$:
@@ 3.1.2 Approximating the Optimal Control Distribution via Kullback-Leibler (KL) Divergence Minimization
$=\min_{\pi}\{\mathbb{E}_{\pi(\tau)}[-\log p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)]$
$\qquad\quad+\mathbb{E}_{\pi(\tau)}\left[\log(\pi(\tau)-\log p(\tau))\right]\},$ (4)
where the final transformation uses Eq. (3). Assuming the optimality likelihood $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$ follows a Boltzmann distribution, Eq. (4) can be marginalized over states as:
@@
are substituted into Eq. (4) [^2] Equation (5) can also be derived via ELBO minimization.. Equation (6) defines the Boltzmann distribution, a probability density function that assigns higher values to trajectories with lower costs. As shown in Fig. 4, the Boltzmann distribution’s shape changes with the temperature parameter $\lambda$: smaller $\lambda$ produces distributions more sharply peaked around the lowest-cost trajectory. While the Boltzmann distribution can be derived from maximum entropy principles, there is no strict theoretical requirement to use it, and alternative distributions are possible [^3] The Cross Entropy Method employs a threshold-based distribution as the optimality likelihood [botev2013cross].. However, the Boltzmann distribution is commonly used due to its practical tractability.
Examining Eq. (5), we see that the optimal control distribution $\pi^{*}$ minimizes a weighted sum (with weight determined by temperature parameter $\lambda$) of the expected cost function and the KL divergence from the prior distribution. Thus, $\pi^{*}(\mathbf{u}_{0:T-1})$ is a distribution that minimizes the expected cost while constraining deviation from the prior distribution.
Figure 4: Cost function and Boltzmann distribution. The Boltzmann distribution assigns higher probability to trajectories with lower costs, with its shape controlled by the temperature parameter $\lambda$. The example cost function is $J(u)=0.6u^{2}\sin(5\pi u)$.
Figure 5: Variation of the optimal control distribution with temperature parameter $\lambda$. The cost function is the same as in Fig. 4. The prior distribution is $p(u)=\mathcal{N}(-2.0,1.0)$.
3.1.3 Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
Finally, we derive Eq. (2) from Eq. (5) using Lagrange multipliers $\alpha$. Specifically, we minimize Eq. (5) subject to the probability normalization constraint $\int\pi(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}=1$. Defining the Lagrangian as
@@ 3.1.3 Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
we obtain:
@@
From the normalization constraint
@@
Therefore, we finally complete the derivation of Eq. (2).

### v2

@@ Abstract
This paper provides a tutorial and a survey of probabilistic inference-based model predictive control (PI-MPC) for robotics. PI-MPC defines an optimal control distribution shaped by trajectory cost, a control prior, and a temperature parameter. In the tutorial part, we derive this view and describe action generation via variational inference, highlighting Model Predictive Path Integral (MPPI) control as a representative algorithm. In the survey part, we organize the PI-MPC literature around the principal design choices identified in the tutorial: prior distribution design, multi-modal distribution handling, constraint satisfaction, scalability to high-degree-of-freedom robots, hardware acceleration, and theoretical foundations. Overall, this paper aims to serve as a coherent entry point for researchers and practitioners interested in understanding, implementing, and extending PI-MPC.
@@
Model Predictive Control, Probabilistic Inference, Sampling-Based MPC, Model Predictive Path Integral Control, Control as Inference
articletype: REVIEW ARTICLE
@@ 1 Introduction
Establishing a universal control theory applicable to diverse robots and tasks has long been a fundamental goal in robotics research. Model Predictive Control (MPC) offers an intuitive and general framework for robot control by optimizing actions while predicting future states over a finite horizon. To apply MPC to real-world robots and tasks, we may need to formulate complex optimal control problems that capture the intricacies of robot dynamics and task requirements. However, these formulations often exhibit strong nonlinearity, making them intractable for traditional numerical optimization methods widely familiar in the control field. In many robotic applications, what we need is not necessarily strong theoretical guarantees (e.g., infinite-horizon stability), but an easy-to-use optimization framework that can reliably solve a broad class of optimal control problems with minimal implementation effort.
@@
Based on the above background, probabilistic inference-based MPC, hereafter referred to as PI-MPC, has gained popularity for many applications recently. As illustrated in Fig. 1, PI-MPC represents the optimal control problem as an optimal control distribution and solves it using variational inference (VI) techniques. This approach belongs to the class of sampling-based MPC methods, which randomly sample robot actions and determine optimal behaviors from these samples. Consequently, it is applicable to optimal control problems with arbitrary dynamics and cost functions.
@@
Figure 1: Overview of probabilistic inference-based model predictive control. The framework consists of two stages: (1) deriving the optimal control distribution from the finite-horizon optimal control problem, and (2) sampling control input sequences from this distribution for execution. This perspective reformulates optimal control as probabilistic inference.
We can trace attempts to solve optimal control problems through probabilistic inference back to the early 2000s, beginning with path integral control [kappen2005path] and formulating stochastic optimal control as Kullback-Leibler (KL) divergence minimization [todorov2008general, kappen2012optimal]. It then extends to information-theoretic frameworks through Model Predictive Path Integral Control (MPPI) [williams2017information, williams2018information], and reaches connections to Reinforcement Learning (RL) through the general framework of Control as Inference [levine2018reinforcement]. Despite its theoretical elegance and growing impact in robotics and decision-making, its formulation is often scattered across different research communities, making it difficult for control theorists and robotic practitioners to obtain a coherent understanding. Moreover, existing tutorials tend to focus either on RL-based formulations [levine2018reinforcement] or on path integral optimal control [kazim2024recent]. While the former is commonly framed in terms of machine learning rather than MPC, the latter is primarily developed within control theory and does not always make its connection to variational or probabilistic inference explicit.
This paper aims to provide a comprehensive tutorial and survey on PI-MPC, serving as a guide for both theorists and practitioners. In the first half, we explain the theoretical foundations of PI-MPC, primarily based on [levine2018reinforcement, williams2018information, wang2021variational, abdolmaleki2018maximum]. Specifically, as illustrated in Fig. 1, we describe the approach in two steps. First, we derive the optimal control distribution from the optimal control problem and discuss its probabilistic interpretation and key characteristics. Second, we explain how to generate actions from the optimal control distribution using VI techniques, taking MPPI as a representative example [^1] We open-source a PyTorch-based MPPI implementation, available at https://github.com/kohonda/mppi_playground. In the latter half, we survey key challenges, representative methods, and recent advances in PI-MPC, including fundamental topics and cutting-edge research directions.
2 Tutorial on Probabilistic Inference-Based Model Predictive Control
2.1 Definition of Optimal Control Problem
@@
Here, $J$ is the cost function; $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ denote the state and control input at time $t$. We write the control sequence as $\mathbf{u}_{0:T-1}=[\mathbf{u}_{0},\dots,\mathbf{u}_{T-1}]$ and the corresponding state sequence as $\mathbf{x}_{0:T}=[\mathbf{x}_{0},\dots,\mathbf{x}_{T}]$. The dynamics are given by $\mathbf{f}$, and $\mathbf{x}_{0}$ is the observed initial state. Most optimal control problems also impose constraints on the state and the control input. In PI-MPC, input constraints can be incorporated into the dynamics, whereas state constraints are typically handled indirectly by embedding them in the cost function (e.g.,, via indicator functions). For details, refer to the experiments in [williams2018information] and Section 3.3.
Given $J$ and $\mathbf{f}$, the controller solves Eq. (1) online and applies the resulting action sequence $\mathbf{u}^{*}_{0:T-1}$. When $J$ and $\mathbf{f}$ admit simple structures (e.g., quadratic objectives and differentiable dynamics), numerical optimization can be efficient. However, when $J$ and/or $\mathbf{f}$ are highly nonlinear or do not provide reliable analytical derivatives, standard optimization methods can become difficult to apply in real time.
@@ 2.2 Motivations for Probabilistic Inference-Based MPC
When numerical optimization proves infeasible, our remaining option is computationally intensive sampling-based methods. The simplest approach is the random shooting method: we sample many candidate control sequences $\mathbf{u}_{0:T-1}$, roll out the resulting trajectories using Eq. (1b), evaluate each trajectory cost via Eq. (1a), and choose the sequence with the lowest cost. As shown in Fig. 2, random shooting is highly sample-inefficient and quickly suffers from the curse of dimensionality: the number of required samples grows exponentially with the input dimension.
Figure 2: Comparison of sampling-based MPC methods in a vehicle obstacle avoidance task [honda2024stein]. Random shooting (Left) selects a single best trajectory from samples, resulting in poor sample efficiency, whereas MPPI (Right) estimates an optimal control distribution and achieves significantly improved efficiency through distribution-level inference.
To overcome the sample efficiency challenges, we therefore introduce a probabilistic inference framework. Specifically, as illustrated in Fig. 1, we approximate the probabilistic distribution of optimal control inputs (hereafter referred to as the optimal control distribution) using samples and generate control input sequences from this distribution. Rather than searching for a single best sequence among random samples, we infer a distribution over good control sequences; this distribution-level view can substantially improve sample efficiency (Fig. 2). Beyond sample efficiency, casting control as distribution optimization offers additional benefits:
  • It naturally represents stochastic variability in behavior.
  • Most computations are embarrassingly parallel, making CPU/GPU parallelization straightforward.
  • The overall pipeline can be implemented end-to-end in a differentiable manner (e.g., in PyTorch).
2.3 Framework of Probabilistic Inference-Based MPC
In this section, we outline the basic PI-MPC framework. As shown in Fig. 1, the scheme involves two key steps: deriving the optimal control distribution from the optimal control problem in Eq. (1) and then generating control inputs to apply to the robot from this optimal control distribution. Thus, two fundamental questions arise:
@@ 2.3 Framework of Probabilistic Inference-Based MPC
2.3.1 Derivation of the Optimal Control Distribution
We begin by deriving the optimal control distribution from the optimal control problem in Eq. (1). We first state the key result: under mild assumptions, the probability density function $\pi^{*}$ of the optimal control distribution takes the following simple form:
@@ 2.3.1 Derivation of the Optimal Control Distribution
Here, $Z$ is the normalization constant, $\lambda$ is the temperature parameter, $J_{\tau}(\mathbf{u}_{0:T-1})$ is the trajectory cost obtained from the control input $\mathbf{u}_{0:T-1}$ via Eq. (1), and $p(\mathbf{u}_{0:T-1})$ represents the prior distribution over control inputs. Thus, the optimal control distribution is proportional to the product of a Boltzmann distribution and a prior distribution [^2] A similar expression appears in Direct Preference Optimization [rafailov2023direct], which has been used for aligning large language models.. The Boltzmann term assigns higher probability to lower-cost trajectories; we will discuss it in more detail later. We now derive this expression in three steps.
Graphical Model Representation of the Optimal Control Problem
Figure 3: Graphical-model representation of the finite-horizon optimal control problem. State transitions depend on current states and control inputs, while optimality variables encode trajectory cost as probabilistic likelihood. This formulation enables variational inference of the optimal control distribution.
We rewrite the optimal control problem in Eq. (1) as a graphical model. In a graphical model, nodes denote random variables and edges encode conditional dependencies. Figure 3 shows the resulting factorization. In particular, the next state $\mathbf{x}_{t+1}$ depends on the current state $\mathbf{x}_{t}$ and control $\mathbf{u}_{t}$, which corresponds to the dynamics in Eq. (1b).
The key additional node is $\mathcal{O}_{t}$, the optimality variable. It is a virtual binary variable indicating whether the state–action pair at time $t$ is “optimal” ($\mathcal{O}_{t}=1$) or not ($\mathcal{O}_{t}=0$), and it will be used to encode the cost function in Eq. (1a) as a likelihood term.
Based on this model, we define the trajectory $\tau=[\mathbf{x}_{0:T},\mathbf{u}_{0:T-1}]$ and the sequence of optimality variables $\mathcal{O}_{0:T}=[\mathcal{O}_{0},\dots,\mathcal{O}_{T}]$. Conditioning on the event that all time steps are optimal, Bayes’ rule yields the posterior trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$:
@@
where $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$ is the optimality likelihood, which assigns higher probability to trajectories that are more “optimal” under the cost.
Approximating the Optimal Control Distribution via Kullback-Leibler Divergence Minimization
We modeled the optimal trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ in Eq. (3). However, since the virtual binary variable $\mathcal{O}_{0:T}$ cannot be computed in advance, this distribution is intractable. Additionally, since the prior distribution $p(\mathbf{u}_{0:T-1})$ included in $p(\tau)$ may contain system-side input noise, it is not necessarily directly controllable by us. We therefore introduce a controllable trajectory distribution $\pi(\tau)$ and use variational inference (VI) to make it close to $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$. Specifically, we minimize the KL divergence $\mathbb{D}_{\rm{KL}}$:
@@
$=\min_{\pi}\{\mathbb{E}_{\pi(\tau)}[-\log p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)]+\mathbb{E}_{\pi(\tau)}\left[\log\pi(\tau)-\log p(\tau)\right]\},$ (4)
where the final transformation uses Eq. (3). If we assume a Boltzmann form for the optimality likelihood $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$, then marginalizing out the states yields:
@@
are substituted into Eq. (4) [^3] Equation (5) can also be derived via ELBO minimization.. Equation (6) defines the Boltzmann distribution, which assigns higher probability to lower-cost trajectories. As illustrated in Fig. 4, the temperature parameter $\lambda$ controls its sharpness: smaller $\lambda$ concentrates mass near the minimum-cost region. Although the Boltzmann form can be motivated by maximum-entropy arguments, it is not the only choice; alternative likelihood models are possible [^4] For example, the Cross-Entropy Method uses a thresholded likelihood [botev2013cross].. We use the Boltzmann form because it leads to a simple and practical inference procedure.
Equation (5) shows that $\pi^{*}$ trades off two terms: the expected cost and the KL divergence to the prior. The temperature $\lambda$ determines the relative weight of these terms. In other words, $\pi^{*}(\mathbf{u}_{0:T-1})$ favors low expected cost while remaining close to the prior distribution.
Figure 4: Relationship between the cost function and the Boltzmann distribution. The Boltzmann distribution assigns higher probability to trajectories with lower costs, and its sharpness is governed by the temperature parameter $\lambda$. Smaller $\lambda$ concentrates probability around the minimum-cost region, whereas larger $\lambda$ yields a smoother distribution. The example cost function is $J(u)=0.6u^{2}+\sin(5\pi u)$ with time horizon $T=1$.
Figure 5: Temperature $\lambda$-dependent variation of the optimal control distribution. The optimal control distribution is formed by the product of the Boltzmann distribution and the prior distribution. Larger temperature values increase the influence of the prior, yielding smoother distributions closer to the prior, whereas smaller temperatures emphasize cost optimality and produce sharper modes. The cost function is the same as in Fig. 4, and the prior distribution is $p(u)=\mathcal{N}(-2.0,1.0)$.
Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
Finally, we derive Eq. (2) from the KL-minimization problem in Eq. (5) using a Lagrange multiplier $\alpha$. We minimize Eq. (5) subject to the normalization constraint $\int\pi(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}=1$. Define the Lagrangian
@@
and set its functional derivative with respect to $\pi$ to zero:
@@
Enforcing the normalization constraint gives
@@
Therefore, we recover Eq. (2).